# Graph-Based Approximate Nearest Neighbor Search Revisited: Theoretical Analysis and Optimization

### v2

@@ 2.2 Probabilistic Framework
Probability Space for SNG construction. We introduce the probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ denotes the sample space containing all possible sequences of nearest neighbor selections and pruning steps during overall SNG construction. The set $\mathcal{F}$ is the corresponding $\sigma$-algebra, capturing all measurable events in $\Omega$ , and $\mathbb{P}$ is the associated probability measure. We further define the natural filtration $\{\mathcal{F}_{t}\}_{t\geq 1}$, representing the information available up to iteration $t$. That is, $\mathcal{F}_{t}$ is generated by the sequence of nearest neighbor selections and candidate sets:
$\mathcal{F}_{t}=\sigma\left((p^{*}_{1},S_{1}),(p^{*}_{2},S_{2}),\dots,(p_{t},S_{t})\right)$
@@
where $p^{*}_{i}=\arg\min_{s\in S_{i}}\|p-s\|$ is the nearest neighbor selected at iteration $i$. The update rule for candidate set is given by: $S_{t+1}=S_{t}\setminus(\{p^{\prime}\in S_{t}:\|p-p^{\prime}\|>\alpha\cdot\|p^{*}-p^{\prime}\|\}\cup\{p^{*}_{t+1}\})$. This formulation models the pruning process as a sequence of progressively revealed decisions. In each iteration, the state of the system is updated based on local geometric information, following a stochastic rule defined by the data distribution and the pruning condition.
@@
We now show that the construction process exhibits a consistent dynamic trend. Intuitively, in each iteration, the number of processed points strictly increases, while the size of the candidate set strictly decreases. This monotonic behavior can be formally captured through the martingale framework.
@@
A stochastic process $\{X_{t}\}_{t\geq 0}$ is a martingale with respect to the natural filtration $\{\mathcal{F}_{t}\}$ if, for all $t\geq 0$:
@@
It is a supermartingale if $\mathbb{E}[X_{t+1}|\mathcal{F}_{t}]\leq X_{t}$, and a submartingale if $\mathbb{E}[X_{t+1}|\mathcal{F}_{t}]\geq X_{t}$.
In the context of SNG construction, the size of candidate set $|S_{t}|$ forms a supermartingale, while the size of the processed set $|S^{\prime}_{t}|$ forms a submartingale, both with respect to the natural filtration $\{\mathcal{F}_{t}\}$. This is consistent with the fact that at each iteration, the candidate set $S_{t}$ can only decrease in expectation in each step, whereas the processed set $S^{\prime}_{t}$ increases monotonically as the algorithm processes. From the SNG pruning rule and the probability framework, we summarize the following properties of processed sets $S^{\prime}_{t}$, which will facilitate the analysis of the construction process. These properties directly follow from our previous results and do not require separate proof.
@@
Given a dataset $P\subset\mathbb{R}^{d}$ consisting of $n$ points, the SNG construction of $p\in P$ satisfies the following properties:
(1) Monotonicity: The processed sets form a nested sequence $\emptyset=S_{0}^{\prime}\subset S^{\prime}_{1}\subset S^{\prime}_{2}\subset\cdots\subset S^{\prime}_{T}$, with a submartingale $|S^{\prime}_{t}|$.
(2) Termination: The final processed set includes n1n-1 points, excluding only the current point $p$. (3) Degree property: Out-degree of $p$ equals the number of iterations, as each iteration adds exactly one directed edge from $p$ to its nearest neighbor in the current candidate set.
@@
2.3 Distribution Assumption
For our theoretical analysis, we assume that large-scale, high-dimensional data points are independently and uniformly distributed. This assumption is also adopted in NSG19 , where it serves as the basis for the theoretical analysis of MSNET. Although real-world datasets may exhibit non-uniformity, our core probabilistic framework remains valid and offers potential for broader generalization. A more detailed discussion of this assumption is provided in Appendix G.
3 Degree Bounds of SNG
We focus on bounding the degree of SNG. As shown in Lemma 4, the out-degree of each point equals the number of pruning iterations during construction. Thus, our analysis shifts to examining the dynamics of the pruning process and estimating the total number of iterations required and the total number of pruning iterations required during index construction.
First, we aim to demonstrate the existence of an initial fast pruning phase that significantly reduces the candidate set within a few iterations. To this end, we begin by characterizing the pruning probability at each step of the SNG construction in the following Lemma 6. This is a key component in quantifying how many points are removed from the candidate set in expectation. The proof is deferred to Appendix C.1. In particular, Appendix C.2 provides a detailed discussion of the applicability of Lemma 6.
Lemma 6 (Pruning Probablity in SNG Construction).
Given a dataset $P\subset\mathbb{R}^{d}$ consisting of $n$ points independently and uniformly distributed within a ball of radius $\rho_{0}$, centered at $p$, we consider the SNG construction processed on the center point $p$ with $\alpha=1$. The probability $\pi_{t}$ that a point $p^{\prime}$ is pruned in the $t$-th iteration is given by $\pi_{t}=\frac{I_{1-(\rho_{t}/\rho_{0})^{2}}\left(\frac{d+1}{2},\frac{1}{2}\right)-(\rho_{t}/\rho_{0})^{d}\cdot I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{2\left(1-(\rho_{t}/\rho_{0})^{d}\right)},$ where $I_{x}(a,b)$ denotes the regularized incomplete Beta function, and $\rho_{t}$ is the distance between the nearest neighbor and $p$ in ttht-th iteration. For large $d$, this probability satisfies $\frac{I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{2}<\pi_{t}<\frac{1}{2}.$
To motivate the study of the fast pruning behavior, we examine a low-dimensional setting using the $M$-$t$ level. As a concrete example, in two dimensions, we show that only four iterations are sufficient to prune over 80% of the candidate points.
@@ 3 Degree Bounds of SNG
Recall that in the SNG construction of $p\in P$, we iteratively prune a candidate set $S_{i}$, starting with $S_{0}=P\setminus{p}$. At iteration $i=1,2,$..., nearest neighbor is selected from $S_{i-1}$, and prune other points with probability $\pi_{i}$. Specifically, the number of points pruned in iteration $i$, denoted by $\Delta S_{i}$, follows a binomial distribution $\Delta S_{i}\sim\text{Bin}(|S_{i-1}|-1,\pi_{i})$, where the subtraction accounts for the exclusion of the selected nearest neighbor from the pruning. After $t$ iterations, $S^{\prime}_{t}$ includes all pruned points and selected nearest neighbors. Thus, we have $|S^{\prime}_{t}|=\sum_{i=1}^{t}(\Delta S_{i}+1)$ for $t=2,3,\dots$ , and for $t=1$, $|S^{\prime}_{1}|=\Delta S_{1}+1$, since each iteration contributes one nearest neighbor and $\Delta S_{i}$ pruned points.
From Lemma 6, we know that for $d=2$, the pruning probability at each iteration satisfies $\pi_{i}>1/3$ (see the end of Appendix C.1 for details). To estimate the pruning progress, we assume a minimal pruning rate of $\pi_{\min}=\frac{1}{3}$ per iteration. Expected number of points pruned in iteration $i$ is $E[\Delta S_{i}]=\pi_{i}(|S_{i-1}|-1)$, and the variance is $\text{Var}(\Delta S_{i})=\pi_{i}(1-\pi_{i})(|S_{i-1}|-1)$. Since $\pi_{i}(1-\pi_{i})<\frac{1}{3}$, we bound the normalized variance $\text{Var}(\frac{\Delta S_{i}}{n-1})=\pi_{i}(1-\pi_{i})(|S_{i-1}|-1)/(n-1)^{2}\leq\frac{1}{3(n-1)}$.
If each iteration prunes at least a fraction $\pi_{\min}^{i}$, where $\pi_{\min}^{1}=1/3$, $\pi_{\min}^{2}=2/9$, $\pi_{\min}^{3}=4/27$, and $\pi_{\min}^{4}=8/81$, the cumulative expected fraction pruned after four iterations is 0.802.
To ensure this holds with high probability, we control the deviation of $\Delta S_{i}$ from its expectation by selecting constants $\epsilon_{i}$, such that:
$\epsilon_{1}<\frac{E[\Delta S_{1}]}{n-1}-\frac{1}{3},\quad\epsilon_{2}<\frac{E[\Delta S_{2}]}{n-1}-\frac{2}{9},\quad\epsilon_{3}<\frac{E[\Delta S_{3}]}{n-1}-\frac{4}{27},\quad\epsilon_{4}<\frac{E[\Delta S_{4}]}{n-1}-\frac{8}{81}.$
By Chebyshev’s inequality, we obtain the following concentration bound for each iteration:
$\Pr\left(\left|\frac{\Delta S_{i}}{n-1}-\frac{p_{i}(|S_{i-1}|-1)}{n-1}\right|\leq\epsilon_{i}\right)\geq 1-\frac{\text{Var}\left(\frac{\Delta S_{i}}{n-1}\right)}{\epsilon_{i}^{2}}\geq 1-\frac{1}{3(n-1)^{2}\epsilon_{i}^{2}}.$
Assuming the deviations remain within $\epsilon_{i}$ for all four iterations, the overall probability that cumulative number of processed points satisfies $|S^{\prime}_{4}|\geq 0.8(n-1)$ is at least
$\Pr(|S^{\prime}_{4}|\geq 0.8(n-1))\geq\prod_{i=1}^{4}\left(1-\frac{1}{3(n-1)^{2}\epsilon_{i}^{2}}\right)=1-o(1).$
@@
In high-dimensional settings, the same intuition is reflected in the concept of sublinear-time progress: a linear number of points is processed within sublinear iterations. This is formalized below.
@@
Given a dataset $P\subset\mathbb{R}^{d}$ with $n$ points distributed uniformly:
(1) For any constants $\nu_{1},\nu_{2}\in(0,1)$ with $\nu_{1}>\nu_{2}$, with probability 1, the SNG construction reaches the $n^{\nu_{1}}$-$O(n^{\nu_{2}})$ level.
(2) For any constant $\nu\in(0,1)$, with probability 1, the SNG construction reaches the $(n-n^{1-\nu})$-$O(n^{\nu})$ level.
Lemma 8 is proven in Appendix C.3, and will be critical in bounding the total number of pruning iterations in Theorem 1. This naturally raises the question of whether the pruning process continues to progress rapidly beyond the initial phase. Empirical results presented in Appendix H.2 demonstrate that, in practice, the pruning process undergoes a significant slowdown after the initial fast pruning phase, often reaching a plateau phase.
To analyze the plateau phase, we observe that the submartingale $S^{\prime}_{t}$ exhibits bounded per-step variance throughout the pruning process. Leveraging this property, we introduce the Differential Equation Method (DEM) (see Lemma 11), a standard tool for approximating the evolution of bounded-variance submartingales. By combining the analyses of both the initial fast pruning phase and the plateau phase, we obtain a comprehensive characterization of the pruning dynamics.
We now revisit our main result, Theorem 1, and provide a sketch of the proof. The full proof is provided in Appendix C.4.

### v3

@@ 3. OPT-SNG Probabilistic Framework Setup
Probability Space for SNG construction. We introduce a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ denotes the sample space containing all possible sequences of nearest neighbor selections and pruning outcomes during the overall SNG construction. The $\sigma$-algebra $\mathcal{F}$ captures all measurable events on $\Omega$ , and $\mathbb{P}$ is the associated probability measure. We further define the natural filtration $\{\mathcal{F}_{t}\}_{t\geq 1}$, representing the information available up to iteration $t$. Specifically,
$\mathcal{F}_{t}=\sigma\left((p^{*}_{1},S_{1}),(p^{*}_{2},S_{2}),\dots,(p_{t}^{*},S_{t})\right)$
@@
where $p^{*}_{i}=\arg\min_{s\in S_{i}}\|p-s\|$ is the nearest neighbor selected at iteration $i$. The candidate set is updated according to the pruning rule $S_{t+1}=S_{t}\setminus(\{p^{\prime}\in S_{t}:\|p-p^{\prime}\|\geq\alpha\cdot\|p^{*}-p^{\prime}\|\}\cup\{p^{*}_{t+1}\})$ which removes both the selected nearest neighbor and all candidates blocked by it. This formulation captures the pruning process as a sequence of progressively revealed decisions.
@@
Martingale Perspective. We now formalize the dynamic trend of the construction process. Intuitively, at each iteration, the processed set strictly grows while the candidate set shrinks. This monotonic behavior can be rigorously characterized using martingale theory.
@@
A stochastic process $\{X_{t}\}_{t\geq 0}$ is a martingale with respect to a filtration $\{\mathcal{F}_{t}\}$ if
@@
for all $t\geq 0$. It is a supermartingale if $\mathbb{E}[X_{t+1}|\mathcal{F}_{t}]\leq X_{t}$, and a submartingale if $\mathbb{E}[X_{t+1}|\mathcal{F}_{t}]\geq X_{t}$.
Table 1. Summary of Notation
Symbol Description
$P$ Dataset of $n$ points in $\mathbb{R}^{d}$
$S_{t}$ Candidate set at iteration $t$
$p^{*}_{t}$ Nearest neighbor of $p$ in $S_{t}$ at iteration $t$
$\mathcal{F}_{t}$ Natural filtration up to iteration $t$
$\Delta S_{t}$ Number of points pruned in iteration $t$
$\pi_{t}$ Pruning probability at iteration $t$
$R$ Truncation parameter (maximum degree limit)
$M$ Number of processed points at a given level
$\alpha$ Relaxation parameter
$t$ Iteration index (time step)
In the context of SNG construction, the size of the candidate set $|S_{t}|$ forms a supermartingale, while the size of the processed set $|S^{\prime}_{t}|$ forms a submartingale with respect to the natural filtration $\{\mathcal{F}_{t}\}$. This reflects the fact that, in expectation, candidates are eliminated while processed points accumulate over time. We focus on $|S^{\prime}_{t}|$ rather than $|S_{t}|$ because, as complementary sets, $|S^{\prime}_{t}|$’s monotonic growth is suited to our analytical framework.
From the SNG pruning rule and the probability framework, we summarize the following properties of processed sets $S^{\prime}_{t}$, which will facilitate the analysis of the construction process. These properties follow directly from our previous results and do not require separate proof.
@@
Given a dataset $P\subset\mathbb{R}^{d}$ consisting of $n$ points, the SNG construction of $p\in P$ satisfies the properties as follows:
  • Monotonicity: The processed sets form a nested sequence $\emptyset=S_{0}^{\prime}\subset S^{\prime}_{1}\subset S^{\prime}_{2}\subset\cdots\subset S^{\prime}_{T}$, with a submartingale $|S^{\prime}_{t}|$.
  • Termination: The final processed set includes n1n-1 points, excluding only the current point $p$.
  • Degree property: Out-degree of $p$ equals the number of iterations, as each iteration adds exactly one directed edge from $p$ to its nearest neighbor in the current candidate set.
Based on this framework, we introduce the following metric to measure the progression of graph construction.
@@
4. Degree Bounds of SNG
In this section, we establish upper bounds on the degree of SNG. As shown in Lemma 3, the out-degree of each point equals the number of pruning iterations during construction, since exactly one edge is added in each iteration. Consequently, bounding the degree reduces to analyzing the dynamics of the pruning process and estimating the total number of iterations required for graph construction. For completeness, all detailed proofs of lemmas and theorems are provided in the full version of this paper on our GitHub repository.
4.1. Two-Phase Behavior: Empirical Motivation
To motivate our theoretical analysis, we first present empirical evidence revealing a distinctive two-phase behavior in SNG construction. Figure 2 illustrates two representative cases from the SIFT1M dataset: a fast case completing in only 31 iterations and a slow case requiring over 1000 iterations. Despite this disparity in total iterations, both cases exhibit a consistent structural pattern: an initial phase of rapid candidate reduction, followed by a prolonged plateau phase during which progress slows significantly.
This empirical observation strongly suggests that a phase-wise analysis is both natural and necessary for deriving degree bounds. In the remainder of this section, we formalize this intuition. Specifically, Lemma 3 establishes rapid progress in the initial phase, which in turn enables a degree bound in Theorem 4.
(a) Fast pruning
(b) Slow pruning
Figure 2. Evolution of processed and candidate sets during SNG construction. (a) shows a fast pruning case terminating in 31 iterations. (b) shows a slow pruning case requiring over 1,000 iterations.
4.2. Pruning Probability Analysis
We begin by analyzing the pruning probability at each iteration, which governs the rate at which candidate points are eliminated. This is a key ingredient for quantifying the speed of the initial fast pruning phase.
Lemma 0 (Pruning Probability in SNG Construction).
Given a dataset $P\subset\mathbb{R}^{d}$ consisting of $n$ points independently and uniformly distributed within a ball of radius $\rho_{0}$, centered at $p$, we consider the SNG construction processed on the center point $p$ with $\alpha=1$. The probability $\pi_{t}$ that a point $p^{\prime}$ is pruned in the $t$-th iteration is given by
$\pi_{t}=\frac{I_{1-(\rho_{t}/\rho_{0})^{2}}\left(\frac{d+1}{2},\frac{1}{2}\right)-(\rho_{t}/\rho_{0})^{d}\cdot I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{2\left(1-(\rho_{t}/\rho_{0})^{d}\right)},$
where $I_{x}(a,b)$ denotes the regularized incomplete Beta function, and $\rho_{t}$ is the distance between the nearest neighbor and $p$ in ttht-th iteration. For large $d$, this probability satisfies
$\frac{I_{0.75}\left(\frac{d+1}{2},\frac{1}{2}\right)}{2}<\pi_{t}<\frac{1}{2}.$
4.3. Two-Dimensional Case
To build intuition, we first analyze a low-dimensional setting using the $M$-$t$ level. In particular, for the two-dimensional case, we show that only four iterations suffice to prune over 80% of all candidate points with high probability.
@@ 4.3. Two-Dimensional Case
Recall that in the SNG construction of $p\in P$, we iteratively prune a candidate set $S_{i}$, starting with $S_{0}=P\setminus{\{p\}}$. At iteration $i=1,2,\ldots$, the nearest neighbor is selected from $S_{i-1}$, and the remaining points are pruned independently with probability $\pi_{i}$.
Let $\Delta S_{i}$ denote the number of points pruned at iteration $i$. Since the selected nearest neighbor is excluded from pruning, we have $\Delta S_{i}\sim\mathrm{Bin}\bigl(|S_{i-1}|-1,\pi_{i}\bigr).$ After $t$ iterations, the processed set $S^{\prime}_{t}$ consists of all pruned points and the selected nearest neighbors. Therefore, $|S^{\prime}_{t}|=\sum_{i=1}^{t}(\Delta S_{i}+1),\ t=2,3,\ldots,$ with $|S^{\prime}_{1}|=\Delta S_{1}+1$, since each iteration contributes exactly one selected nearest neighbor in addition to the pruned points.
From Lemma 1, we know that for $d=2$, the pruning probability at each iteration satisfies $\pi_{i}>1/3$. To estimate the pruning progress, we conservatively assume a uniform lower bound $\pi_{\min}=1/3$ across all iterations.
The expected number of pruned points at iteration $i$ is $\mathbb{E}[\Delta S_{i}]=\pi_{i}(|S_{i-1}|-1),$ and the variance is $\mathrm{Var}(\Delta S_{i})=\pi_{i}(1-\pi_{i})(|S_{i-1}|-1).$ Since $\pi_{i}(1-\pi_{i})<1/3$, the normalized variance can be bounded as
(1) $\mathrm{Var}\left(\frac{\Delta S_{i}}{n-1}\right)=\frac{\pi_{i}(1-\pi_{i})(|S_{i-1}|-1)}{(n-1)^{2}}\leq\frac{1}{3(n-1)}.$
Under the minimal pruning rate assumption, the expected fraction of points pruned at each iteration is at least
$\pi_{\min}^{1}=\tfrac{1}{3},\quad\pi_{\min}^{2}=\tfrac{2}{9},\quad\pi_{\min}^{3}=\tfrac{4}{27},\quad\pi_{\min}^{4}=\tfrac{8}{81}.$
Consequently, the cumulative expected fraction of processed points after four iterations exceeds $0.8$.
To ensure that this behavior holds with high probability, we control the deviation of $\Delta S_{i}$ from its expectation by choosing constants $\epsilon_{i}>0$ such that
$\epsilon_{1}<\frac{\mathbb{E}[\Delta S_{1}]}{n-1}-\frac{1}{3},\quad\epsilon_{2}<\frac{\mathbb{E}[\Delta S_{2}]}{n-1}-\frac{2}{9},$
$\epsilon_{3}<\frac{\mathbb{E}[\Delta S_{3}]}{n-1}-\frac{4}{27},\quad\epsilon_{4}<\frac{\mathbb{E}[\Delta S_{4}]}{n-1}-\frac{8}{81}.$
Applying Chebyshev’s inequality yields, for each iteration $i$,
(2) $\Pr\left(\left|\frac{\Delta S_{i}}{n-1}-\frac{\pi_{i}(|S_{i-1}|-1)}{n-1}\right|\leq\epsilon_{i}\right)$ $\geq 1-\frac{\mathrm{Var}\left(\frac{\Delta S_{i}}{n-1}\right)}{\epsilon_{i}^{2}}$
(3) $\geq 1-\frac{1}{3(n-1)^{2}\epsilon_{i}^{2}}.$
Assuming that all four deviations remain within their respective bounds, the probability that the cumulative number of processed points satisfies $|S^{\prime}_{4}|\geq 0.8(n-1)$ is bounded below by
(4) $\Pr(|S^{\prime}_{4}|\geq 0.8(n-1))\geq\prod_{i=1}^{4}\left(1-\frac{1}{3(n-1)^{2}\epsilon_{i}^{2}}\right)=1-o(1).$
@@
In high-dimensional settings, the same phenomenon manifests as sublinear-time progress: a linear number of points is processed within sublinear iterations. This behavior is formalized in the following lemma.
@@
Given a dataset $P\subset\mathbb{R}^{d}$ with $n$ points distributed uniformly, for all relaxation parameter $\alpha>0$:
(1) For any constants $\nu_{1},\nu_{2}\in(0,1)$ with $\nu_{1}>\nu_{2}$, with probability one, the SNG construction reaches the $n^{\nu_{1}}$-$O(n^{\nu_{2}})$ level.
(2) For any constant $\nu\in(0,1)$, with probability one, the SNG construction reaches the $(n-n^{1-\nu})$-$O(n^{\nu})$ level.
Proof.
Part (1). Fix constants $\nu_{1}>\nu_{2}$ and define the first-passage time $t_{1}:=\inf\{t\geq 0:|S^{\prime}_{t}|\geq n^{\nu_{1}}\}.$ We prove that $t_{1}=O(n^{\nu_{2}})$ almost surely, i.e., there exists a constant $C>0$ such that $\Pr(t_{1}\leq Cn^{\nu_{2}})=1.$ Equivalently, it suffices to show that for any fixed $C>0$,
(5) $\Pr\!\left(|S^{\prime}_{\lfloor Cn^{\nu_{2}}\rfloor}|<n^{\nu_{1}}\ \text{i.o.}\right)=0,$
where "i.o.” (infinitely often) denotes the event
(6) $\limsup_{n\to\infty}\left\{|S^{\prime}_{\lfloor Cn^{\nu_{2}}\rfloor}|<n^{\nu_{1}}\right\}=\bigcap_{n_{0}=1}^{\infty}\ \bigcup_{n\geq n_{0}}\left\{|S^{\prime}_{\lfloor Cn^{\nu_{2}}\rfloor}|<n^{\nu_{1}}\right\}.$
Let $t=\lfloor Cn^{\nu_{2}}\rfloor$. Since $|S^{\prime}_{t}|=\sum_{i=1}^{t}(\Delta S_{i}+1),$ we obtain the deterministic lower bound $|S^{\prime}_{t}|\;\geq\;t\cdot\min_{1\leq i\leq t}(\Delta S_{i}+1),$ which implies
(7) $\Pr(|S^{\prime}_{t}|<n^{\nu_{1}})\;\leq\;\Pr\!\left(\min_{1\leq i\leq t}\Delta S_{i}<\frac{n^{\nu_{1}}-t}{t}\right).$
Since $\nu_{1}>\nu_{2}$, for sufficiently large $n$ we have $n^{\nu_{1}}>2t$, and hence
(8) $\frac{n^{\nu_{1}}-t}{t}\;\geq\;\frac{1}{2C}\,n^{\nu_{1}-\nu_{2}}.$
Recall that $\Delta S_{i}\sim\mathrm{Bin}(|S_{i-1}|-1,\pi_{i})$ with $\mathbb{E}[\Delta S_{i}]=\pi_{i}(|S_{i-1}|-1).$ Moreover, the sampling probability admits the uniform lower bound $\pi_{i}\geq\tfrac{1}{2}I_{0.75}\!\left(\tfrac{d+1}{2},\tfrac{1}{2}\right).$ Applying the Chernoff bound (Chernoff), for any $\delta\in(0,1)$,
(9) $\Pr\!\big(\Delta S_{i}\leq(1-\delta)\mathbb{E}[\Delta S_{i}]\big)\;$ $\leq\;\exp\!\Big(-\tfrac{\delta^{2}}{2}\mathbb{E}[\Delta S_{i}]\Big)$
(10) $\leq\;\exp\!\Big(-\tfrac{\delta^{2}}{4}\big(|S_{i-1}|-1\big)\,I_{0.75}\!\left(\tfrac{d+1}{2},\tfrac{1}{2}\right)\Big).$
For all $i\leq t$ such that $|S_{i-1}|\geq n/2$, we have $\mathbb{E}[\Delta S_{i}]\geq c_{0}n$ for some constant $c_{0}>0$. Choosing $\delta\in(0,1)$ such that
(11) $(1-\delta)\mathbb{E}[\Delta S_{i}]\;\geq\;\frac{1}{2C}\,n^{\nu_{1}-\nu_{2}},$
we obtain $\Pr(|S^{\prime}_{t}|<n^{\nu_{1}})\;\leq\;\exp(-cn)$ for some constant $c>0$ and all sufficiently large $n$. Consequently,
(12) $\sum_{n=1}^{\infty}\Pr\!\left(|S^{\prime}_{\lfloor Cn^{\nu_{2}}\rfloor}|<n^{\nu_{1}}\right)<\infty.$
By the Borel–Cantelli lemma (see Lemma 1 in Appendix),
(13) $\Pr\!\left(|S^{\prime}_{\lfloor Cn^{\nu_{2}}\rfloor}|<n^{\nu_{1}}\ \text{i.o.}\right)=0,$
which completes the proof of Part (i).
Part (2). Let $\nu\in(2/3,1)$ and define $t_{2}:=\inf\{t\geq 0:\ |S^{\prime}_{t}|\geq n-n^{1-\nu}\}.$ We show that $t_{2}=O(n^{\nu})$ almost surely. Fix $C>0$ and set $t=\lfloor Cn^{\nu}\rfloor$. As above,
(14) $\Pr\!\left(|S^{\prime}_{t}|<n-n^{1-\nu}\right)\;\leq\;\Pr\!\left(\min_{1\leq i\leq t}\Delta S_{i}<\frac{n-n^{1-\nu}-t}{t}\right).$
For sufficiently large $n$, we have $t\leq n/2$ and thus
(15) $\frac{n-n^{1-\nu}-t}{t}\;\geq\;\frac{1}{2C}\,n^{1-\nu}.$
Using the same Chernoff bound and the same uniform lower bound on $\pi_{i}$, we again obtain an exponentially small upper bound
(16) $\Pr\!\left(|S^{\prime}_{t}|<n-n^{1-\nu}\right)\;\leq\;\exp(-c^{\prime}n)$
for some constant $c^{\prime}>0$. Therefore,
(17) $\sum_{n=1}^{\infty}\Pr\!\left(|S^{\prime}_{\lfloor Cn^{\nu}\rfloor}|<n-n^{1-\nu}\right)<\infty.$
Applying the Borel–Cantelli lemma yields
(18) $\Pr\!\left(|S^{\prime}_{\lfloor Cn^{\nu}\rfloor}|<n-n^{1-\nu}\ \text{i.o.}\right)=0,$
and hence $\Pr(t_{2}=O(n^{\nu}))=1$.
4.4. Main Degree Bound
Lemma 3 establishes that the initial phase rapidly processes a large fraction of points. However, prior empirical observations suggest that this rapid progress does not persist throughout the entire construction. The pruning process undergoes a significant slowdown after the initial fast pruning phase, often reaching a plateau phase.
To analyze this plateau phase, we exploit a key fact: the submartingale ${|S^{\prime}_{t}|}_{t\geq 0}$ has bounded per-step variance along the pruning trajectory. This enables us to apply the Differential Equation Method (DEM) (DEM1; DEM2; DEM3), a standard technique for approximating bounded-variance stochastic processes by deterministic trajectories. DEM yields an ordinary differential equation that captures the expected evolution of $|S^{\prime}_{t}|$ in the plateau regime. Combining the initial fast-pruning bound from Lemma 3 with the DEM-based characterization of the plateau phase gives a unified view of the pruning dynamics. We now sketch the proof of our main result, Theorem 4. Detailed proofs of all lemmas and theorems are provided in the full version of this work on our GitHub repository.